OpenStudy (anonymous):

solve each system of equations: 3) x+2y=0 4x-z= 4 5y+z=-1

myininaya (myininaya):

do you want to do it by matrix?

OpenStudy (anonymous):

no by substitution or elimination

myininaya (myininaya):

ok well if we add the second two equations we get 4x+5y=3 so we also have x+2y=0 solve this one for x and get x=-2y plugged this into the other equation and solve for y

myininaya (myininaya):

4(-2y)+5y=3 <solve for y

myininaya (myininaya):

after you get y remember that x=-2y so finding x will be like eating cake

myininaya (myininaya):

hopefully you get the same z for both of those bottom equations

OpenStudy (anonymous):

x+2y=0 x=-2y 4x-z=4 4(-2y)-z=4 -8y-z=4 -8y-4=z 5y+z = -1 5y + (-8y-4) = -1 5y -8y-4 = -1 -3y-4=-1 -3y=3 y=-1

OpenStudy (anonymous):

x+2y=0 x+2(-1) = 0 x-2 = 0 x=2 4x-z=4 4(2)-z=4 8-z=4 -z=-4 z=4

myininaya (myininaya):

-3y=3 => y=-1 x=-2y=-2(-1)=2 --------- 4(2)-z=4 => z=4 5(-1)+z=-1 => z=4 ok good just checking

OpenStudy (anonymous):

oh ok thanks! haha i always mess up on these :/